### v1

@@ 3.4.3 Imaging Diagnosis (Generative QA)
The imaging diagnosis task requires LLMs to understand medical images and use detailed textual descriptions to identify key features in findings, such as tumors, inflammation, or other pathological changes, providing imaging diagnostic reports $R$, i.e., $R=\text{LLM}_{\textit{image\_diag}}(K)$. This task evaluates the LLM’s ability to analyze and interpret medical imaging reports for diagnostic support.
@@ 4.1 Models
To comprehensively analyze strengths and weaknesses of a broad selection of competitive LLMs on different tasks and departments in ClinicalBench, we evaluate 11 general LLMs, including GPT-4 [23], GPT-3.5 [23], Gemini-Pro-1.0 [11], InternLM2-20B-Chat [6], Yi-34B-Chat [39], Qwen-72B-Chat [2], Baichuan2-13B-Chat [38], Claude3 [1], ChatGLM3-6B [10], BlueLM-7B-Chat [32], and Spark-3 [13]. We also evaluate 6 Chinese medical LLMs, including HuatuoGPT2-34B [7], WiNGPT2-14B-Chat [27], PULSE-20B [41], Taiyi-LLM [21], DISC-MedLLM[3], and BianQue-2 [8]. To ensure the reproducibility of the experiments, we set the temperature parameter to 00 or set the do_sample parameter to $False$. The experimental evaluation is completed through API calls using 8 NVIDIA A6000 GPUs.
@@ 4.2 Evaluation Metrics (ClinicalMetrics)
We propose four novel metrics to precisely measure the effectiveness of LLMs in department guide and their clinical diagnostic capabilities, as follows. Additionally, we use metrics such as accuracy, BLEU [25], ROUGE [19], and BertScore [40] to evaluate the experimental results. See Appendix  A.1. for details.
Department Win Rate (DWR) measures the relative performance of LLMs in clinical diagnosis across different medical departments, which is calculated as Eq. 1,
$\text{DWR}=\frac{1}{m}\sum_{j=1}^{m}\left(\frac{1}{n}\sum_{i=1}^{n}\left(n-(r_{i}^{j}-1)\right)\right)$ (1)
@@
Department Instruction Following Rate (DIFR) measures the degree to which LLMs follow task instructions when generating the quantity and names of ranked departments. It is the average of two sub-metrics: DIFR-Q and DIFR-N.
Department Quantity Following Rate (DIFR-Q) measures the consistency between the quantity of top $K^{\prime}$ departments generated by the model and the quantity of $K$ departments specified in the task instructions, which is calculated as Eq. 2,
[…]
@@
Department Name Following Rate (DIFR-N) measures the consistency of the department names generated by the model with the names in a predefined list of 24 departments, which is calculated as Eq. 3,
[…]
@@
Comprehensive Diagnostic Accuracy (CDR) measures the accuracy of LLMs in simultaneously predicting correct department guide and disease diagnosis, which is calculated as Eq. 4,
$\text{CDR}=\frac{\sum_{i=1}^{N}\text{S}_{guide}(i)\times\text{S}_{diagnosis}(i)}{N}$ (4)
where $\text{S}_{guide}(i)$ and $\text{S}_{diagnosis}(i)$ are indicator functions of correctness for the $i$-th sample in the guide and diagnosis tasks, respectively. If the prediction is correct, the function value is 1, and 0 otherwise. $N$ is the total number of test samples.
@@
$\text{Acceptability}=\left(\frac{1}{N}\sum_{i=1}^{N}\text{CDR}(i)\right)\times\left(\frac{1}{NKM}\sum_{i=1}^{N}\sum_{k=1}^{K}\sum_{m=1}^{M}\text{Metric}_{m}^{k}(i)\right)$ (5)
@@ 4.3 Results & Analysis
Table 2 shows the performance of 17 LLMs on ClinicalBench. The prompt templates for various tasks can be found in Appendix Table 10-12. Note that all experiments in this work are conducted on the original Chinese samples of ClinicalBench [^4] We also provide the code to support evaluations using the English translations of ClinicalBench..
@@
Overall Performance. InternLM2 achieves the top Overall DWR of 14.91 and Avg. 51.98, indicating its superiority over other models across multiple departments with balanced performance. We hypothesize that the diverse pre-training data, the thorough data filtering methods, and careful executions of pre-training, supervised fine-tuning, and alignments through a novel reinforcement learning [6] may have contributed to the superior performance of InternLM2. GPT-4 performs the best in simultaneously predicting the correct department and disease diagnosis (CDR), showing its ability to complete 33.27% of end-to-end medical consultations. Moreover, GPT-4 demonstrates the effectiveness of its content quality and accuracy with a top acceptability score of 11.37.
Overall Findings. From the task dimension, we observe that existing LLMs are not yet fully capable of handling end-to-end clinical diagnosis tasks. Interestingly, we find that the clinical diagnostic capabilities of general LLMs are better than those of the specialized medical LLMs; hence, how to improve specialized models with medical data remains an open challenge. From the department dimension, Figure 4 shows the rankings of different models across various departments, demonstrating that
no single LLM excels in all departmental domains. This finding aligns well with the reality of human medical experts, where no single expert can master all departmental domains.
@@ 5.1 ClinicalAgent: Methodology
Departmental Guide: The patient $PA$ presents a chief complaint $CC$ to the patient navigator agent $PN$. Based on $CC$ and the prior knowledge $PK$ of LLM rankings obtained in Table 2, agent $PN$ arranges $N$ clinicians from the $K$ most relevant departments in the consultation room to prepare a preliminary consultation for $PA$. The purpose of this stage is to quickly and accurately identify the patient’s initial needs and arrange for the appropriate specialty medical team to diagnose.
Preliminary Consultation: The patient $PA$ meets multiple clinicians from different departments in the consultation room, such as pediatrics clinician powered by GPT-4, orthopedic clinician powered by InternLM2. Each clinician makes a preliminary diagnosis $PD$ based on the $CC$, medical history $H$, and physical examination $V$, and decide on further laboratory $LE$ and imageological $IE$ examinations.
Laboratory Examination: According to the recommendations $LE$ from the preliminary consultation, the patient $PA$ proceeds to the biochemical testing room for necessary laboratory examination results, such as blood tests and urine analysis. The output is the corresponding laboratory results $A$.
Imageological Examination: According to the recommendations $IE$ from the preliminary consultation, the patient $PA$ undergoes necessary imageological tests, such as X-rays, CT scans, or MRI. The output is the corresponding medical imaging report $R$.
Final Consultation: The patient $PA$ returns to the consultation room and provides the $A$ and $R$ to the team from the preliminary consultation. Based on the newly obtained laboratory and imaging reference information, the medical team provides the final diagnosis $FD$ and treatment plans $TP$.
Medical Treatment: The patient $PA$ receives corresponding treatment based on $FD$ and $TP$, including treatments conducted in the treatment room and medication dispensed in the pharmacy. Finally, the patient leaves the hospital after completing the treatment.
@@ 5.2 ClinicalAgent: Evaluations
Table 4 and Figure 6 illustrate the performance of 4 LLMs and 3 configurations of ClinicalAgent using the three different evaluation methods. Tables 15-36 in the Appendix provide detailed case studies, including case information, diagnostic results from human doctors and seven models. Using both Automatic Evaluation and Human Evaluation methods, ClinicalAgent (Agent#K@N) based on department scheduling and doctor allocation strategies leads top-performing LLMs with a large margin. With GPT-4o evaluation, GPT-4 achieves the highest score of 90.93, which is probably attributable to the evaluation model’s preference for its own generated answers [42]. Even so, ClinicalAgent (Agent#K@N) receives competitive GPT-4o scores up to 89.40. Overall, ClinicalAgent (Agent#K@N) excel in handling complex diagnosis tasks, showcasing the effectiveness of collaborative diagnostic strategies across multiple departments and doctors.
Interestingly, different configurations of ClinicalAgent exhibit varying performance; overall, Agent#1@3 performs the best. We conduct deeper analysis of these variants. Figure 6 shows that Agent#1@3 excels in fluency, relevance, and proficiency in medicine, benefiting from integrating diagnostic opinions from multiple doctors within the same department, which effectively fills in gaps and reduces misdiagnosis. However, due to its focus on a single department, Agent#1@3 falls slightly short in completeness and is more suitable for clinical diagnosis of a single disease or multiple closely related diseases. Figure 6 shows that Agent#3@1 excels in comprehensiveness and proficiency in medicine, but its fluency and relevance are slightly compromised. This may be attributed to errors in Department Guide Acc@3, leading to noisy opinions from doctors from irrelevant department. Overall, Agent#3@1 benefits from diverse expertise across departments and generates high-quality diagnostic results more suitable for joint clinical diagnosis of multiple complex or loosely related diseases. In contrast, while Agent#1@1’s performance is inferior to the collaboration of multiple experts, it still significantly outperforms other top-performing LLMs, demonstrating the effectiveness of flexible and optimized department scheduling and doctor allocation strategies in clinical diagnostic agents.
@@ 7 Limitations
To date, no medical agents have been evaluated on real clinical data, including the recent Agent Hospital [18]. Existing agents are primarily designed to handle multiple-choice question answering and are not suitable for simulating complex real-world scenarios. Therefore, ClinicalAgent cannot be directly compared with other agents. However, as a foundational method, ClinicalAgent can be well integrated with other agents.
@@ Appendix A Appendix
A.1 Evaluation Metrics
@@
Accuracy indicates the proportion of samples correctly classified by the department to the total number of samples. BLEU [25] measures the degree of overlap between the predicted output and the reference output, ranging from 1-gram to 4-gram. Rouge [19] primarily evaluates based on the longest common subsequence found between the generated texts and the human-written reference ones. BertScore [40] leverages a pre-trained BERT model and its deep contextual representations to gauge the semantic similarity between the generated and reference texts.
A.2 ClinicalAgent Algorithm
Departmental Guide: The patient $PA$ presents a simple chief complaint $CC$ to the patient navigator agent $PN$ powered by $\text{LLM}_{Yi}$, where $\text{LLM}_{Yi}$ donates to the LLM Yi-Chat [39]. Based on $CC$ and the prior knowledge $PK$ obtained in Table 2, agent $PN$ arranges $N$ clinicians $CL_{j}(j=1,2,...,N)$ from the $K$ most relevant departments $DE_{i}(i=1,2,...,K)$ in the consultation room to prepare a preliminary consultation for $PA$, where $K$ denotes the number of departments to be scheduled, and $N$ denotes the number of the top clinicians assigned to each department. This process is represented as $\{DE,CL\}=\bigcup_{i=1}^{K}\left(DE_{i},\{CL_{j}\}_{j=1}^{N}\right),\text{ where }DE_{i}\in PN_{LLM_{Yi}}(CC),\{CL_{j}\}\in PK(DE_{i})$. The purpose of this stage is to quickly and accurately identify the patient’s initial needs and arrange for the appropriate specialty medical team to diagnose.
Preliminary Consultation: The patient $PA$ meets with multiple clinicians from different departments in the consultation room, such as pediatrics clinician $CL_{GPT-4}^{Pediatrics}$ powered by GPT-4, orthopedic clinician $CL_{InternLM2}^{Orthopedics}$ powered by InternLM2, and hematology clinician $CL_{Gemini-Pro}^{Hematology}$ powered by Gemini-Pro. Each clinician makes a preliminary diagnosis $PD$ based on the chief complaint $CC$, medical history $H$, and physical examination $V$, and decide on further laboratory $LE$ and imageological $IE$ examinations. This process is represented as $\{PD,LE,IE\}=CL_{Top1}^{DE_{Top1}}\left(\bigcup_{i=1}^{K}\bigcup_{j=1}^{N}CL_{j}^{DE_{i}}(CC,H,V)\right)$, where $CL_{Top1}^{DE_{Top1}}$ is the top doctor in the department most related to $CC$.
Laboratory Examination: According to the recommendations $LE$ from the preliminary consultation, the patient $PA$ proceeds to the biochemical testing room for necessary laboratory examination results, such as blood tests and urine analysis. The output is the corresponding laboratory results $A$. This process is represented as $A=\bigcup_{{test}\in LE}Lab(test,PA)$.
Imageological Examination: According to the recommendations $IE$ from the preliminary consultation, the patient $PA$ undergoes necessary imageological tests, such as X-rays, CT scans, or MRI. The output is the corresponding medical imaging report $R$. This process is represented as $R=\bigcup_{test\in IE}Img(test,PA)$.
Final Consultation: The patient $PA$ returns to the consultation room and provides the laboratory results $A$ and imageological report $R$ to the team from the preliminary consultation. Based on the newly obtained laboratory and imaging data, the medical team provides the final diagnosis $FD$ and treatment plans $TP$. This process is represented as $(FD,TP)=CL_{Top1}^{DE_{Top1}}\left(\bigcup_{i=1}^{K}\bigcup_{j=1}^{N}CL_{j}^{DE_{i}}(PD,CC,H,V,A,R)\right)$.
Medical Treatment: The patient $PA$ receives corresponding treatment based on $FD$ and $TP$, including treatments conducted in the treatment room and medication dispensed in the pharmacy. Finally, the patient leaves the hospital after completing the treatment.

### v2

@@ 3.4.3 Imaging Diagnosis (Generative QA)
The Imaging Diagnosis (ID) task requires LLMs to understand detailed textual reports of medical images, identify key features of lesions, such as tumors, inflammation, or other pathological changes, and provide imaging diagnosis results $\mathcal{IDR}$, i.e., $\mathcal{IDR}=\text{ID}(\mathcal{IR})$. This task evaluates the LLM’s ability to analyze and interpret medical imaging reports for diagnostic support.
@@ 4.1 Models
To comprehensively analyze strengths and weaknesses of a broad selection of competitive LLMs on different tasks and departments in ClinicalBench, we evaluate 11 general LLMs, including GPT-4 [23], GPT-3.5 [23], Gemini-Pro-1.0 [11], InternLM2-20B-Chat [6], Yi-34B-Chat [37], Qwen-72B-Chat [2], Baichuan2-13B-Chat [36], Claude3 [1], ChatGLM3-6B [10], BlueLM-7B-Chat [32], and Spark-3 [13]. We also evaluate 6 Chinese medical LLMs, including HuatuoGPT2-34B [7], WiNGPT2-14B-Chat [27], PULSE-20B [39], Taiyi-LLM [21], DISC-MedLLM[3], and BianQue-2 [8]. To ensure the reproducibility of the experiments, we use greedy decoding for each LLM (set the temperature parameter to 00 or set the do_sample parameter to False). The experimental evaluation is completed through API calls and 8 NVIDIA A6000 GPUs.
@@ 4.2 Evaluation Metrics (ClinicalMetrics)
We propose four novel metrics to precisely measure the effectiveness of LLMs in department guide and their clinical diagnostic capabilities, as follows. Additionally, we use metrics such as accuracy, BLEU [25], ROUGE [19], and BertScore [38] to evaluate the experimental results. For more details and the handling of medical synonyms, please refer to Appendix A.2.
Department Win Rate ($\mathsf{DWR}$) measures the relative performance of LLMs in clinical diagnosis across different medical departments, which is calculated as Eq. 1,
$\mathsf{DWR}=\frac{1}{m}\sum_{j=1}^{m}\left(\frac{1}{n}\sum_{i=1}^{n}\left(n-(\mathsf{r}_{i}^{j}-1)\right)\right)$ (1)
@@
Department Instruction Following Rate ($\mathsf{DIFR}$) measures the degree to which LLMs follow task instructions when generating the quantity and names of ranked departments. It is the average of two sub-metrics: $\mathsf{DIFR\text{-}Q}$ and $\mathsf{DIFR\text{-}N}$.
Department Quantity Following Rate ($\mathsf{DIFR\text{-}Q}$) measures the consistency between the quantity of top $K^{\prime}$ departments generated by the model and the quantity of $K$ departments specified in the task instructions, which is calculated as Eq. 2,
$\mathsf{DIFR\text{-}Q}=\frac{\sum_{i=1}^{N}\mathbf{S}_{\{K^{\prime}_{i}=K\}}}{N}$ (2)
@@
Department Name Following Rate ($\mathsf{DIFR\text{-}N}$) measures the consistency of the department names generated by the model with the names in a predefined list of 24 departments, which is calculated as Eq. 3,
$\mathsf{DIFR\text{-}N}=\frac{\sum_{i=1}^{N}\sum_{j=1}^{K^{\prime}_{i}}\text{match}(i,j)}{\sum_{i=1}^{N}K^{\prime}_{i}}$ (3)
@@
Comprehensive Diagnostic Rate ($\mathsf{CDR}$) measures the accuracy of LLMs in simultaneously predicting correct department guide and disease diagnosis, which is calculated as Eq. 4,
$\mathsf{CDR}=\frac{\sum_{i=1}^{N}\text{S}_{\text{guide}}(i)\times\text{S}_{\text{diagnosis}}(i)}{N}$ (4)
where $\text{S}_{\text{guide}}(i)$ and $\text{S}_{\text{diagnosis}}(i)$ are indicator functions of correctness for the $i$-th sample in the guide and diagnosis tasks, respectively. If the prediction is correct, the function value is 1, and 0 otherwise. $N$ is the total number of test samples.
@@
$\mathsf{Acceptability}=\left(\frac{1}{N}\sum_{i=1}^{N}\mathsf{CDR}(i)\right)\times\left(\frac{1}{NKM}\sum_{i=1}^{N}\sum_{k=1}^{K}\sum_{m=1}^{M}\text{Metric}_{m}^{k}(i)\right)$ (5)
@@ 4.3 Results & Analysis
Table 2 shows the performance of 17 LLMs on ClinicalBench. Note that all experiments in this work are conducted on the original Chinese samples of ClinicalBench [^4] Due to high costs, we are unable to conduct these repetitive experiments, but we provide code to support the evaluation using the English version of ClinicalBench..
@@
Overall Performance. InternLM2 achieves the top $\mathsf{DWR}$ of 14.91 and Avg. 51.98, indicating its superiority over other models across multiple departments with balanced performance. We hypothesize that the diverse pre-training data, thorough data filtering methods, and careful executions of pre-training, supervised fine-tuning, and alignments through a novel reinforcement learning [6] may have contributed to the superior performance of InternLM2. GPT-4 performs the best in simultaneously predicting the correct department and disease diagnosis ($\mathsf{CDR}$), showing its ability to complete 33.27% of end-to-end medical consultations. Moreover, GPT-4 demonstrates the effectiveness of its content quality and accuracy with a top acceptability score of 11.37.
Overall Findings From the task dimension, we observe that existing LLMs are not yet fully capable of handling end-to-end clinical diagnosis tasks. Interestingly, we find that the clinical diagnostic capabilities of general LLMs are better than those of the specialized medical LLMs; hence, how to improve specialized models with medical data remains an open challenge. From the department dimension, Figure 4 shows the rankings of different models across various departments, demonstrating that no single LLM excels in all departmental domains. This finding aligns well with the reality of human medical experts, where no single expert can master all departmental domains.
@@ 5.1 ClinicalAgent: Methodology
Departmental Guide: The patient $\mathcal{PA}$ presents a chief complaint $\mathcal{CC}$ to the patient navigator agent $\mathcal{PN}$. Based on $\mathcal{CC}$ and the prior knowledge $\mathcal{PK}$ of LLM rankings obtained in Table 2, agent $\mathcal{PN}$ arranges $\mathcal{N}$ clinicians from the $\mathcal{K}$ most relevant departments in the consultation room to prepare a preliminary consultation for $\mathcal{PA}$. The purpose of this stage is to quickly and accurately identify the patient’s initial needs and arrange for the appropriate specialty medical team to diagnose.
Preliminary Consultation: The $\mathcal{PA}$ consults with multiple clinicians from different departments in the consultation room, such as pediatrics clinician powered by GPT-4, orthopedic clinician powered by InternLM2, and hematology clinician powered by Gemini-Pro. Each clinician makes a preliminary diagnosis $\mathcal{P}$ based on the $\mathcal{CC}$, $\mathcal{MH}$, and $\mathcal{PE}$, and decide on further laboratory tests and imageological examinations.
Laboratory Examination: According to the preliminary consultation advice, the $\mathcal{PA}$ proceeds to the biochemistry laboratory for a series of necessary examinations, such as blood tests and urine analysis. Subsequently, based on the laboratory report $\mathcal{LR}$ generated by the medical instruments, the biochemist provides the laboratory diagnostic results $\mathcal{LDR}$.
Imageological Examination: According to the preliminary consultation advice, the $\mathcal{PA}$ undergoes necessary imageological examinations in the radiology room, such as X-rays, CT scans, or MRI. Subsequently, the radiologist provides an imaging diagnosis result $\mathcal{IDR}$ based on the radiological reports $\mathcal{IR}$ of these medical images.
Final Consultation: The $\mathcal{PA}$ returns to the consultation room and provides $\mathcal{LDR}$ and $\mathcal{IDR}$ to the medical team. Based on the newly acquired reference information, the medical team conducts a step-by-step analysis and ultimately provides results including $\mathcal{B}$, $\mathcal{D}$, $\mathcal{F}$, $\mathcal{G}$, and $\mathcal{T}$.
Medical Treatment: The $\mathcal{PA}$ receives corresponding treatment based on $\mathcal{G}$ and $\mathcal{T}$, including treatments conducted in the treatment room and medication dispensed in the pharmacy. Finally, the patient leaves the hospital after completing the treatment.
@@ 5.2 ClinicalAgent: Evaluations
Table 4 and Figure 6 illustrate the performance of 4 LLMs and 3 configurations of ClinicalAgent using the three different evaluation methods. Tables 10-31 in the Appendix provide detailed case studies, including case information, diagnostic results from human doctors and seven models. Using both Automatic Evaluation and Human Evaluation methods, ClinicalAgent (Agent#K@N) based on department scheduling and doctor allocation strategies leads top-performing LLMs with a large margin. With GPT-4o evaluation, GPT-4 achieves the highest score of 90.93, which is probably attributable to the evaluation model’s preference for its own generated answers [41]. Even so, ClinicalAgent (Agent#K@N) receives competitive GPT-4o scores up to 89.40. Overall, ClinicalAgent (Agent#K@N) excel in handling complex diagnosis tasks, showcasing the effectiveness of collaborative diagnostic strategies across multiple departments and doctors.
@@
Overall Findings Interestingly, different configurations of ClinicalAgent exhibit varying performance. Overall, Agent#1@3 performs the best. We conduct deeper analysis of these variants. Figure 6 shows that Agent#1@3 excels in fluency, relevance, and proficiency in medicine, benefiting from integrating diagnostic opinions from multiple doctors within the same department, which effectively fills in gaps and reduces misdiagnosis. However, due to its focus on a single department, Agent#1@3 falls slightly short in completeness and is more suitable for clinical diagnosis of a single disease or multiple closely related diseases. Figure 6 shows that Agent#3@1 excels in comprehensiveness and proficiency in medicine, but its fluency and relevance are slightly compromised. This may be attributed to errors in Department Guide Acc@3, leading to noisy opinions from doctors from irrelevant department. Overall, Agent#3@1 benefits from diverse expertise across departments and generates high-quality diagnostic results more suitable for joint clinical diagnosis of multiple complex or loosely related diseases. In comparison, Agent#1@1, as an ablation study of Agent#3@1 and Agent#1@3, while not performing as well as the collaboration of multiple experts, still clearly outperforms other top-performing LLMs, demonstrating the effectiveness of flexible and optimized department scheduling and doctor allocation strategies in clinical diagnostic agents.
@@ 7 Limitations
To date, no medical agents have been evaluated on real clinical data, including the recent MedAgents [31] and Agent Hospital [18]. Existing agents are primarily designed for handling multiple-choice questions and are not suitable for simulating complex real-world scenarios. Specifically, after carefully reviewing the original paper and code of MedAgents, we find that evaluating it would require significant modifications to its experimental design and prompts. These changes would force MedAgents’ core process, which is designed for multiple-choice questions, to shift toward a generative question-answering approach, likely impacting its performance negatively. In addition to the previously mentioned differences in design goals, data incompatibility, and variations in evaluation metrics and scope, we are particularly concerned that such comparisons may lead to an unfair evaluation of MedAgents. Therefore, ClinicalAgent cannot be directly compared with other agents. However, as a foundational approach, ClinicalAgent can integrate well with other agents.
@@ 8 Ethical Statement
We can provide supporting documents and certification from medical institutions and notary organizations to demonstrate the legality and ethical approval of our data collection process. Since the review process is double-blind, we do not include these documents in the supplementary materials. If you need to review them, please contact us to obtain the relevant materials.
@@ Appendix A Appendix
A.1 The Secure Distribution of ClinicalBench
To further strengthen data security, we adopt an application-based data distribution system. Applicants must review and agree to the terms and commit to maintaining confidentiality. Additionally, we utilize an automated algorithm to embed personalized watermarks in each dataset, based on the applicant’s information. This enables us to swiftly trace the source in the event of a data breach. These measures ensure that ClinicalBench can provide valuable data to the medical AI research community while safeguarding privacy.
As of September 30, 2024, we have successfully distributed ClinicalBench datasets 63 times to researchers from various countries and regions, including the United States, China, the United Kingdom, Germany, and Singapore. This reflects the positive impact ClinicalBench is already making in the AI research community and its contribution to global advancements in medical AI. We are confident that ClinicalBench will continue to play a pivotal role in driving progress in medical AI research.
A.2 Evaluation Metrics
@@
Accuracy indicates the proportion of samples correctly classified by the department to the total number of samples. BLEU [25] measures the degree of overlap between the predicted output and the reference output, ranging from 1-gram to 4-gram. Rouge [19] primarily evaluates based on the longest common subsequence found between the generated texts and the human-written reference ones. BertScore [38] leverages a pre-trained BERT model and its deep contextual representations to gauge the semantic similarity between the generated and reference texts. In the experiment, we use the “bert-base-chinese” [^5] https://huggingface.co/google-bert/bert-base-chinese model along with the “jieba” [^6] https://pypi.org/project/jieba/ Chinese segmentation tool to ensure accurate word tokenization.
To ensure the accuracy of BLEU, ROUGE, and BertScore when evaluating medical synonyms, we implement several key steps. First, we compile a list of synonyms for 150 diseases from medical textbooks and online encyclopedias, encompassing a range of expression variations. To address discrepancies in terminology arising from differences in model training data, we also request that 17 models independently generate synonym lists for each disease. We then merge the synonyms collected from textbooks and encyclopedias with those generated by the models. This combined list is rigorously screened by three medical experts among the authors to produce a standardized synonym list.
Before calculating BLEU, ROUGE, and BertScore, we use this standardized list to align the model-generated terms with the ground truth, minimizing errors due to synonym variations and ensuring the accuracy of the metrics. Furthermore, when distributing the dataset, we include the synonym list to enable other researchers to replicate and validate our work.
A.3 ClinicalAgent Algorithm